The heat of vaporization of CH_2CI_2 is 28.0 kJ/mo
Answer #1

1. With heat of vaporization of CH2Cl2 = 28 kJ/mol, dS for 42 g CH2Cl2 at 39.8 oC would be,

B. 44 J/K

Detailed :

dS = dH/T

moles = 42 g/84 g/mol = 0.5 mol

dH = 28 x 0.5 = 14 kJ

T = 39.8 + 273 = 312.8 K

feed values,

dS = dH/T = 42 x 28/84 x 312.8 = 44 J/K